OpenStudy (anonymous):

Prove that if gcd(a,b)=1, then gcd(a,bc)=gcd(a,c) (Hint: let gcd(a,c)=d and let gcd(a,bc)=e)

OpenStudy (anonymous):

god is it late. i wonder if we can do this without making a mess

OpenStudy (anonymous):

start with \[ax+by=1\]

OpenStudy (anonymous):

then \[ax_1+cy_2=d\] \[ax_2+bcy_2=e\]

OpenStudy (anonymous):

and we want to show that \[d|e,e|d\]

OpenStudy (anonymous):

multiply \[ax+by=1\] by d to get \[adx+bdy=d\]

OpenStudy (anonymous):

no i don't think it is that easy

OpenStudy (anonymous):

but i am on the way i think, although at any time i can screw this up

OpenStudy (anonymous):

good night. i am going to try a couple more lines of this. i know this is the right idea

OpenStudy (anonymous):

it doesn't say anywhere that a b or c are integers so I don't think it is that easy

OpenStudy (anonymous):

we have \[adx+bdy=d=ax_1+cy_1\]

myininaya (myininaya):

i think we are suppose to assume they are integers

OpenStudy (anonymous):

oh of course everything in sight is an integer

OpenStudy (anonymous):

otherwise question doesn't even make sense

OpenStudy (anonymous):

You're right it doesn't make sense if they aren't integers

OpenStudy (anonymous):

oki think i am almost there

OpenStudy (anonymous):

is this line ok \[adx+bdy=d=ax_1+cy_1\]

OpenStudy (anonymous):

it just says \[d=d\]

OpenStudy (anonymous):

That looks right

OpenStudy (anonymous):

damn got stuck again hold on

OpenStudy (anonymous):

ok i think we write \[adx+b(ax_1+cy_1)y=d\] and then \[adx+abx_1y+bcy_1y=d\] and finally \[a(dx+bx_1y)+bcy_1y=d\]

OpenStudy (anonymous):

there got it. and what does this algebra mean?

OpenStudy (anonymous):

the last line is a linear combination of \[a\] and \[bc\] and since \[e=\gcd(a,bc)\] it divides any linear combination of the two

OpenStudy (anonymous):

and therefore \[e|d\]

OpenStudy (anonymous):

i believe it is all there and actually correct. but we are not done. we just showed that e divides d. the next job is to show that d divides e, but that works with the same idea

OpenStudy (anonymous):

again start with \[ax+by=1\] multiply by \[e=ax_2+bcy_2\] to get \[aex+bex=e=ax_2+bcy_2\] and the replace e by \[ax_2+bcy_2\] in that equation and so on.

OpenStudy (anonymous):

You'll have to give me a minute to understand this, but I think it looks right Why can't we just say that acx + bcy = c so a(cx) + bc(y) = c and a(cx) + c(by) = c so a(cx) + bc(y) = a(cx) + c(by) which means that gcd(a, bc) = gcd(a, c) Does this make any sense?

OpenStudy (anonymous):

i have to say that i am too tired to get this but i am almost certain it is wrong . but i cannot give a good reason at the moment.

OpenStudy (anonymous):

I think you are probably right. Thank you for your help!